OpenStudy (anonymous):

Liquid sodium is being considered as an engine coolant. How many grams of liquid sodium (minimum) are needed to absorb 1.00 MJ of energy (in the form of heat) if the temperature of the sodium is not to increase by more than 10.0 °C? Use CP = 30.8 J/(K·mol) for Na(l) at 500 K.

OpenStudy (anonymous):

@zbay

OpenStudy (anonymous):

So what formula do you think we should use here?

OpenStudy (anonymous):

the heat formula?

OpenStudy (anonymous):

q=m x c x delta T ?

OpenStudy (anonymous):

seems to me that one should work I would plug in the numbers insuring the units cancel and solf for m

OpenStudy (anonymous):

i tried to plugged it in but im trying to figure out which one is the initial and which is final, cuz to find delta T you gotta take Final - Initial

OpenStudy (anonymous):

just use 10 C as delta t

OpenStudy (anonymous):

oh, so what about that 500 K ?

OpenStudy (anonymous):

283 k

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok use 283 kalvin the 500k is for the C of sodium at 500k

OpenStudy (anonymous):

wait what? im lost -.-

OpenStudy (anonymous):

so i set up the equation like this: 1.00x10^6 = (m)(30.8)(10) <--- is that right?

OpenStudy (anonymous):

change the 10c to 283k so you go from celcius to kalvin and your units will cancel. I always put the units in so you can make sure they cancel out and you got the answer you are really looking for. you will also have to get your megajoule and joules into the same unit

OpenStudy (anonymous):

Look at your specific heat and get all your other units to match that one

OpenStudy (anonymous):

so this heat equation measured in Kelvin instead of Celcius ?

OpenStudy (anonymous):

so i think it's 1x10^12 = (m)(30.8)(283) <--- it should be alright so solve now

OpenStudy (anonymous):

your specific heat given is measured in Kelvin and not celcius so

OpenStudy (anonymous):

i got such a big number for mass, I'm not even sure if its correct....it's 114726263.1

OpenStudy (anonymous):

alright 1mj is 1x10^6 j and this will leave you in moles so you have to convert that to grams

OpenStudy (anonymous):

wait a minute, i clicked on the "Hint" and it show me this: "A temperature change, ΔT, has the same value in kelvins as it does in degrees Celsius. Thus, ΔT = 10.0 °C = 10.0 K. "

OpenStudy (anonymous):

so it's not 283 but 10 ? o_0

OpenStudy (anonymous):

no q=1,000,000 jolues

OpenStudy (anonymous):

yes, so you change 1x10^6 MJ to J = 1x10^12

OpenStudy (anonymous):

no the system needs to absorb 1 mj and 1 mj = 1x10^6 j

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

can you please put the whole equation together, so i can see better cuz im really confused on which number to put... @.@

OpenStudy (anonymous):

\[1x10^6 j = m (30.8\frac{j}{k*mol})(283.15k)\]

OpenStudy (anonymous):

notice that the kelvins cancell right off the bat and you are left with j/mol

OpenStudy (anonymous):

how to get the mol?

OpenStudy (anonymous):

so after you solve for m what unit is your answer in?

OpenStudy (anonymous):

what do you mean how do you get the mol?

OpenStudy (anonymous):

unit should be in gram

OpenStudy (anonymous):

you don't know how to convert from moles to grams?

OpenStudy (anonymous):

do you have to in this case?

OpenStudy (anonymous):

yes if they want the answer in grams you must convert

OpenStudy (anonymous):

\[114.67 mol Na (\frac{22.99 g Na}{1 mol Na})=342.9 g Na\]

OpenStudy (anonymous):

it's wrong...

OpenStudy (anonymous):

i think the delta T in this case is 10 and not 283

OpenStudy (anonymous):

and if i use that then i would get 3246 mol of Na

OpenStudy (anonymous):

then to grams so...74350 g <-----big number o_0

OpenStudy (anonymous):

\[114.67 mol Na(\frac{22.99 g Na}{1 mol Na})=2636.26 g Na\]

OpenStudy (anonymous):

i messed up on the conversion

OpenStudy (anonymous):

hmm still wrong..

OpenStudy (anonymous):

it sait (Delta T) in K equals to Celcius...do not add 273

OpenStudy (anonymous):

it said*

OpenStudy (anonymous):

my answer was correct.. 74350 g lol =.=

OpenStudy (anonymous):

alright